A drop of water takes x amt. of time to travel the Mississippi River. Determine time.
Drop of water: .5 to 4 millimeters.
Length of river: 2,350 miles.
River travels at 1.2 mph.
Not sure how to solve.

Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52943)

.
A drop of water takes x amt. of time to travel the Mississippi River. Determine time.
Drop of water: .5 to 4 millimeters.
Length of river: 2,350 miles.
River travels at 1.2 mph.
Not sure how to solve.
~~~~~~~~~~~~~~~~


To find time, divide the distance by the average speed.


In the set of given values, find the distance; then find the average speed.


Ignore all other given data, because it is placed in this joke problem with only one single goal to confuse you.


=================


comment from student:   Check my calculations, please:

            2350 / 1.2 = 1958.3333 = 1958 hrs.     1958 / 24 = 81.58333 = 82 days.



My response :   Your calculations are correct.

                        My congratulations (!)



Answer by Alan3354(69443)

A drop of water takes x amt. of time to travel the Mississippi River. Determine time.
Drop of water: .5 to 4 millimeters.
Length of river: 2,350 miles.
River travels at 1.2 mph.
====================================
The size of the drop is not relevant.
----
d = r*t
t = d/r = 2350/1.2 hours (= x hours)